Find f(-2) when f(x)=x2+4x-6. The x is squared.
Found 2 solutions by checkley77, Theo:
Answer by checkley77(12844)

f(x)=x^2+4x-6.
F(-2)=(-2)^2+4*-2-6
F(-2)=4-8-6
F(-2)=-10 ANS.

Answer by Theo(13342)

you just replace x with -2 and solve.

f(x) = x^2 + 4x - 6 is your original equation.

f(-2) = (-2)^2 + 4*(-2) - 6 which you solve to get:

f(-2) = 4 - 8 + 6 = 2